# Usable XAI: 10 Strategies Towards Exploiting Explainability in the LLM Era

### v1

@@ 11.3.1 Uncontrollable Credibility of Emulation
While LLMs can assist in rationale collection and explanation evaluation, their behaviors of collected results may not always match human annotators, primarily due to hallucinated responses in their unfamiliar domains (Ji et al., 2023). This issue leads to unreliable annotations or feedback, as LLMs confidently generate factually incorrect conclusions. The quality of data gathered from this process is compromised, impacting the development of XAI systems. To improve the quality of annotations and feedback, future research could focus on incorporating hallucination detection (Dhuliawala et al., 2023) and retrieval augmented generation (Ren et al., 2023) techniques. These methods could enhance the reliability of LLM outputs, making them more comparable to human-generated content in the context of XAI development.
11.3.2 Ethical Considerations in LLM Annotation
@@
12 Discussion and Conclusion
XAI research is undergoing a significant transformation and experiencing rapid expansion in the era of large models. In previous sections, we have introduced XAI methodologies with an emphasis on their usability. In this final section, we provide a high-level overview of the overarching challenges that persist in the field and suggest directions for future endeavors.
  • Circumvent the interpretability-accuracy tradeoff. The advent of modern LLMs (e.g., ChatGPT) has a significant impact on this tradeoff. Traditionally, in many applications, people are willing to sacrifice a certain degree of performance for better transparency. A corresponding XAI strategy is to train and deploy an inherently interpretable model that mimics the black-box model (Che et al., 2016). However, applying this strategy to LLMs presents a challenge due to the difficulty in identifying an interpretable model that can match the performance levels of LLMs. This requires the creation of XAI strategies that can circumvent this tradeoff, where enhanced interpretability can contribute to improved accuracy. This is consistent with the goal of Usable XAI discussed in this paper.
  • Data-driven AI vs. XAI. Data-driven AI refers to developing AI models that operate based on large volumes of training data. This approach often leads to “black-box” models, as it emphasizes results over the clarity of decision-making pathways. Currently, the development of XAI techniques lags behind the advancement of LLMs because the latter easily scale up with data-driven methods – they ingest gigantic amounts of texts from the Internet to train. However, we believe that XAI might still catch up because of several opportunities. (1) We may run out of data. It was predicted that “we will have exhausted the stock of high-quality language data before 2026” ***https://www.livemint.com/mint-top-newsletter/techtalk20102023.html. Should the accumulation of more data cease to yield substantial improvements, the focus might shift towards enhancing model interpretability to leverage existing data more effectively. (2) The model is relatively stable. As the Transformer architecture of LLMs is pretty mature and stable, it will attract more attention to interpret their inner workings. (3) Leveraging LLMs for XAI. The advancement of XAI research can be accelerated if it can properly use the knowledge and human-like capabilities of LLMs.
  • The objective matters for explanation. During the transition from the era of classical machine learning (when SVMs and decision trees dominate) to the deep learning era (when convolutional and recurrent neural networks became popular), XAI techniques put a strong emphasis on achieving complete transparency within models, as if “anything less than fully transparent is not transparent”. However, as LLMs begin to match or even exceed human capabilities across various tasks, the importance of certain XAI problems shifts. For example, when Recurrent Neural Networks (RNNs) were widely adopted for text generation, we are interested in how the output is linguistically derived because RNNs often produce nonsensical sentences. Nowadays we are less interested in this for LLMs, as they are proficient in generating coherent text. Nevertheless, our focus may shift to explaining how LLMs construct output with factual information, as LLMs are prone to producing hallucinations. Similar observations exist in human cognition, which can be categorized into system-1 and system-2 styles: system-1 handles intuitive and unconscious tasks that are less explainable, while system-2 encompasses logical thinking, planning, and reasoning (Goyal & Bengio, 2022). Given the vast scale and complexity of LLMs, achieving absolute transparency across all aspects of these models appears increasingly unfeasible in the immediate future. Thus, prioritizing meaningful and feasible objectives of explanation, customized for specific tasks, becomes essential in enhancing the utility of AI systems in practical applications.
  • Evaluation remains challenging for XAI in LLMs. Traditional XAI has developed a comprehensive taxonomy of explanation problems and formats, accompanied by clear definitions for each category (Han et al., 2022; Doshi-Velez & Kim, 2017; Rudin et al., 2022). However, the established taxonomy cannot be simply grafted into the study of LLM because of two reasons. First, certain XAI challenges lose their prominence in the context of LLMs, while some approaches become too complex for practical application. Second, while XAI is becoming a common pathway of solving problems for LLMs, the exploration of LLMs’ inner mechanics has branched into various directions. For instance, there has been a notable trend towards leveraging insights from human behavior and limitations to interpret LLMs, such as whether LLMs can lie (Azaria & Mitchell, 2023), can LLMs keep secrets (Mireshghallah et al., 2023), the impact of politeness in prompts on LLMs (Yin et al., 2024b), and even how they can be "hypnotized" (Li et al., 2023e). These diverse approaches have not converged to unified methodologies in interpreting LLM behaviors, which makes it challenging for evaluation. A potential risk is the resultant explanations might give users a false sense that they accurately understand the model, especially when users attempt to shoehorn certain human knowledge or concepts to explain LLMs (Schut et al., 2023).
Conclusion.
@@
In this paper, we hope to guide readers through a crucial yet frequently underappreciated aspect of Explainable AI (XAI) – usability. To this end, we present 10 strategies for advancing Usable XAI within the LLM paradigm, including (1) leveraging explanations to reciprocally enhance LLMs and general AI systems, and (2) enriching XAI approaches by integrating LLM capabilities. Unlocking the potential of XAI’s usability can help address various challenges in LLM such as human alignment. We also provide case studies to several critical topics, aiming to provide resources for interested developers. We further discuss open challenges at the end of each strategy, suggesting directions for future work in this evolving area.
Acknowledgement
The work is, in part, supported by NSF (#IIS-2223768, #IIS-2223769, #IIS-2310261, #DRL-2101104). The views and conclusions in this paper are those of the authors and should not be interpreted as representing any funding agencies.

### v2

@@ 11.3.1. Uncontrollable Credibility of Emulation
While LLMs could assist in rationale collection and explanation evaluation, their provided results can be biased and may not always match human annotators, due to hallucinated responses in their unfamiliar domains (Ji et al., 2023) and that LLMs are still immature (Harding et al., 2023). The quality of data gathered from this process is compromised, impacting the development of XAI systems. To improve the quality of annotations and feedback, future research could focus on incorporating hallucination detection (Dhuliawala et al., 2023) and retrieval augmented generation (Ren et al., 2023) techniques. These methods could enhance the reliability of LLM outputs, making them more comparable to human-generated content in the context of XAI development.
11.3.2. Ethical Considerations in LLM Annotation
@@
12. Conclusion
@@
In this paper, we discuss a crucial yet frequently underappreciated aspect of Explainable AI (XAI) – usability. To this end, we present 10 strategies for advancing Usable XAI within the LLM paradigm, including (1) leveraging explanations to reciprocally enhance LLMs and general AI systems, and (2) enriching XAI approaches by integrating LLM capabilities. Unlocking the potential of XAI’s usability can help address various challenges in LLM such as human alignment. We also provide case studies to several critical topics, aiming to provide resources for interested developers. We further discuss open challenges at the end of each strategy, suggesting directions for future work in this evolving area.